OpenStudy (anonymous):

HELP simplify please!!!! ((x+y)/(x^2 y^3 ))^(1/2)

OpenStudy (anonymous):

(x^2 y^3)^(1/2) = (x^2)^(1/2) * (y^3)^(1/2) = x * y^(3/2) so the entire thing becomes x/(x*y^(3/2)) + y/(x*y^(3/2)) = y^(-3/2) + y^(2/2)/(x*y^(3/2)) = y^(-3/2) + x*y^(-1/2) = y^(-3/2) [ 1 + xy ]

OpenStudy (anonymous):

Nice work. But putting something to the power of -(3/2) is not really 'simplifying' for an algebra course.

OpenStudy (anonymous):

\[(x +y)/\sqrt{x ^{2}y ^{3}}\]\[(x +y)/(xy \sqrt{y})\]If you want to go further\[\sqrt{y}(x +y)/(xy ^{2})\]

OpenStudy (anonymous):

i don't understand any of this, can you explain it more in detail, please

OpenStudy (anonymous):

OK, Jamesm did a good job, but I think the work is too advanced for the course you are taking. In my example, there is not much that can be done with x+y. The denominator\[(x ^{2}y ^{3})^{1/2}\]is the same thing as\[\sqrt{x ^{2}y ^{3}}\]You can simplify the denominator to\[xy \sqrt{y}\]For simplicity, leave it here, go no further\[(x +y)/(xy \sqrt{y})\]

OpenStudy (anonymous):

ok